OpenStudy (anonymous):

If a spherical ball is enlarged so that its surface area is 9 times greater than its original surface area, then the original radius was multiplied by _________.

OpenStudy (anonymous):

tips?

OpenStudy (kinggeorge):

Surface area is proportional to the square of the radius, so if the surface area is 9 times...

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

so

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

we need to find the area of the radius, right?

OpenStudy (callisto):

No...... How can you find the area of radius.....

OpenStudy (anonymous):

true

OpenStudy (kinggeorge):

Area of the radius? lolwut?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

I meant square of the radius

OpenStudy (anonymous):

3 times 3, 1 times 1, 2 times 2, that type of area

OpenStudy (kinggeorge):

I'm not sure if I get what you're saying. Suppose \(A\) is the area of the first sphere, and \(A'\) the area of the second sphere. Likewise, \(r,r'\) are the radii for the first and second sphere respectively. Then\[\frac{A'}{A}=\left(\frac{r'}{r}\right)^2\]You have that \[\frac{A'}{A}=9\]So that means that \[\frac{r'}{r}=...?\]

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

the surface area of a sphere is 4/3 pi cubed, right?

OpenStudy (anonymous):

j

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ehh

OpenStudy (kinggeorge):

\(4\pi r^2\), but we don't need that.

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

well

OpenStudy (anonymous):

lets said that the original surface area was 81

OpenStudy (anonymous):

when you multiply it by 9, it becomes 729

OpenStudy (anonymous):

the radius obviously got bigger

OpenStudy (kinggeorge):

If we let \(x=\frac{r'}{r}\), then can you more easily solve for \(x\) in this equation? \[9=x^2?\]

OpenStudy (anonymous):

3=x

OpenStudy (anonymous):

so the radius got three times bigger?

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

r to the ninth/ r to the first

OpenStudy (anonymous):

dimensions

OpenStudy (anonymous):

r to the third/ r to the first

OpenStudy (kinggeorge):

Substituting back in, and solving, we get \[3r=r'\]So yes, it got 3 times bigger.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

next problem

OpenStudy (callisto):

For similar figures, \[(\frac{s_1}{s_2})^2 = \frac{A_1}{A_2}\]\[(\frac{s_1}{s_2})^3 = \frac{V_1}{V_2}\]